### v3

# Bayesian variable selection in sample selection models using spike-and-slab priors

Abstract
Sample selection models represent a common methodology for correcting bias induced by data missing not at random. These models are not empirically identifiable without exclusion restrictions. In other words, some variables predictive of missingness do not affect the outcome model of interest. The drive to establish this requirement often leads to the inclusion of irrelevant variables in the model. A recent proposal uses adaptive LASSO to circumvent this problem, but its performance depends on the so-called covariance assumption, which can be violated in small to moderate samples. Additionally, there are no tools yet for post-selection inference for this model. To address these challenges, we propose two families of spike-and-slab priors to conduct Bayesian variable selection in sample selection models. These prior structures allow for constructing a Gibbs sampler with tractable conditionals, which is scalable to the dimensions of practical interest. We illustrate the performance of the proposed methodology through a simulation study and present a comparison against adaptive LASSO and stepwise selection. We also provide two applications using publicly available real data. An implementation and code to reproduce the results in this paper can be found at https://github.com/adam-iqbal/selection-spike-slab.
Keywords Gibbs sampling  $\cdot$ Heckman correction  $\cdot$ missing data  $\cdot$ prior elicitation  $\cdot$ scale mixtures
1 Introduction
Sample selection occurs when a portion of the sample is missing non-randomly, leading to a sample that is not representative of the population of interest. This is prevalent in medical and social sciences, where the outcome of interest is non-randomly selected (Heckman, 1976, 1979; Wooldridge, 2010). As a result, the sample becomes biased, potentially distorting the findings and generalizability of the results. Heckman (1976, 1979) introduced a regression model to correct for sample selection bias in the case where the outcome variable is continuous. The main idea is to formulate a two-equation regression model, one for the outcome process and one for the selection process, allowing for correlated errors under the assumption of bivariate normality. This kind of models are known as sample selection models (or Heckman correction). Heckman (1979) proposed a two-stage estimator of the parameters of sample selection models, and established consistency and asymptotic normality of such estimator. After this influential development, several extensions have been proposed to alleviate challenges posed by the assumption of normality of the errors. These include van Hasselt (2011), who extended the outcome model to mixture-of-Gaussian error distributions. Marchenko and Genton (2012), who extended Heckman’s model by modelling the errors in the two-equation model using a bivariate-$t$ distribution, which accounts for heavy tails, coupled with maximum likelihood estimation of the parameters. Ogundimu and Hutton (2016) assumed bivariate skew-normal errors, which account for departures from symmetry. Other extensions include the analysis of binary outcomes subject to sample selection bias. One of the challenges in these models is that the user must specify the variables that enter the outcome and selection models. Entering the same variables in both the outcome and selection submodels often leads to collinearity problems. It has also been shown that even with “exclusion restriction” rules - a covariate included in the selection but not the outcome - collinearity issues may appear in two-step estimators (Leung and Yu, 2000). This challenge points to the need for formal tools for specifying the variables that enter the outcome and selection models.
Recently, the use of variable selection methods in the context of sample selection models has attracted considerable attention. For instance, Ogundimu (2022) used an Adaptive LASSO penalty to select the variables that enter the outcome and selection models. The distributional regression sample selection modelling introduced by Wiemann et al. (2022) allows for the use of Bayesian variable selection priors (shrinkage and spike-and-slab priors). However, Wiemann et al. (2022) only provide an implementation of their proposed model using a Metropolis within Gibbs sampler, which is computationally more expensive than a closed-form Gibbs sampler due to the need for iterative accept-reject Metropolis steps, and requires careful calibration of the priors for the additive components to ensure convergence. In a similar vein, Vera (2023) considered the use of shrinkage and spike-and-slab priors, but they restricted variable selection to only the selection equation, and their implementation similarly relies on a Metropolis within Gibbs sampler.
We propose two families of continuous spike-and-slab priors (George and McCulloch, 1993; Tadesse and Vannucci, 2021) to conduct Bayesian variable selection in sample selection models. The first class of priors allows for the use of scale mixtures of normals for the spike and slab components, which include Laplace and Student-$t$ distributions. The second one is defined conditionally on the marginal variance of the outcome equation, allowing for easier calibration, and also includes scale mixtures of normals for the spike and slab components. We provide a detailed discussion on the calibration of both priors, taking into account the different nature of the outcome and selection models. We derive a closed-form Gibbs sampler for the posterior distribution of the parameters, under both classes of priors, which scales well to the dimensions of interest in practical applications. Posterior samples can be used to jointly select variables for both the selection and outcome equations, without the need for an exclusion restriction, and to produce post-selection inference on the parameters.
The remainder of this paper is organized as follows. In Section 2, we present a brief review of sample selection models and previous literature on Bayesian inference for these models. In Section 3, we present the proposed families of spike-and-slab priors and discuss the calibration of these priors. Section 4 presents details about the Gibbs sampler for the proposed prior structures. Section 5 presents a simulation study aiming at illustrating the performance of the proposed spike-and-slab priors for variable selection. We compare the performance of the proposed methodology with that of adaptive LASSO and forward selection. Indeed, a byproduct of this work is the implementation of stepwise variable selection methods for sample selection models. Section 6 presents real data applications using two popular data sets in the context of sample selection models and a sensitivity analysis for the hyperparameters. We conclude with a discussion and possible extensions in Section 7.
2 Sample selection models
Sample selection models with continuous outcomes consist of two equations, one for the outcome and one for the selection processes, as follows,
[…]
where $\bm{x}_{i}=(x_{i,1},\ldots,x_{i,p})^{\top}\in{\mathbb{R}}^{p}$ and $\bm{w}_{i}=(w_{i,1},\ldots,w_{i,q})^{\top}\in{\mathbb{R}}^{q}$, $i=1,\ldots,n$, are the corresponding covariates, $\bm{\beta}=(\beta_{1},\ldots,\beta_{p})^{\top}\in\mathbb{R}^{p}$ and $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{q})^{\top}\in\mathbb{R}^{q}$ are the regression coefficients, and $\beta_{0}$ and $\alpha_{0}$ are the intercepts. The errors are typically assumed to be distributed according to a bivariate normal distribution
$\begin{pmatrix}\epsilon_{1,i}\\ \epsilon_{2,i}\end{pmatrix}\sim N\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}\sigma^{2}&\sigma\rho\\ \sigma\rho&1\end{pmatrix}\right)\,,$ (2)
where $\sigma>0$ and $\rho\in(-1,1)$. Note that, in equation (2), the variance of the selection equation is set to $1$. This is because we only observe the sign $s_{i}$, so that this parameter cannot be identified. Additionally, the data is only partially observed. That is, we only observe the indicator $s_{i}$ and the selected sample $y_{i}$, where $s_{i}=\mathbb{I}(s^{*}_{i}>0)$ and
$y_{i}=\begin{cases}y_{i}^{*}&\text{if }s_{i}=1\,,\\ \text{missing }&\text{if }s_{i}=0\,.\end{cases}$ (3)
The log-likelihood of the parameters for this model is given by (Toomet and Henningsen, 2008)
$\ell(\alpha_{0},\beta_{0},\bm{\alpha},\bm{\beta},\sigma,\rho)$ $=$ $\sum_{\{s_{i}=0\}}\log\Phi(-\alpha_{0}-\bm{w}_{i}^{\top}\bm{\alpha})$
$+$ $\sum_{\{s_{i}=1\}}\Bigg{(}\log\Phi\left(\frac{\alpha_{0}+\bm{w}_{i}^{\top}\bm{\alpha}+\rho(y_{i}-\beta_{0}-\bm{x}_{i}^{\top}\bm{\beta})/\sigma}{\sqrt{1-\rho^{2}}}\right)$
$\qquad\quad-\frac{1}{2}\left(\frac{y_{i}-\beta_{0}-\bm{x}_{i}^{\top}\bm{\beta}}{\sigma}\right)^{2}-\log(\sigma)-\frac{1}{2}\log(2\pi)\Bigg{)}\,.$
Maximum likelihood estimators of the parameters can be found by maximizing the log-likelihood function using any general-purpose optimization method. An alternative is the two-step method, which uses that
$\mathbb{E}\left(y_{i}\mid\bm{x}_{i},\bm{w}_{i},s_{i}=1\right)=\beta_{0}+\bm{x}_{i}^{\top}\bm{\beta}+\rho\sigma\lambda\left(\alpha_{0}+\bm{w}_{i}^{\top}\bm{\alpha}\right),$ (4)
where $\lambda(t)=\phi(t)/\Phi(t)$ is the Inverse Mills’ ratio, with $\phi(t)$ denoting the standard normal density function and $\Phi(t)$ the standard normal cumulative distribution function. The two-step method first estimates $\alpha_{0}$ and $\bm{\alpha}$ from the selection equation as a probit model, then uses the estimated $\bm{\alpha}$ to fit a linear regression model to (4), treating $\lambda(\alpha_{0}+\bm{w}_{i}^{\top}\bm{\alpha})$ as an additional covariate and $\rho\sigma$ as its coefficient. See Toomet and Henningsen (2008) for more details on frequentist methods.
In the Bayesian framework, van Hasselt (2011) showed that it is possible to construct a tractable Gibbs sampler for the model defined by equations (1)-(3) by using the reparametrization $\tilde{\sigma}^{2}=\sigma^{2}(1-\rho^{2})$ and $\tilde{\rho}=\rho\sigma$. The conditional distributions for $(s^{*}_{i},y^{*}_{i})$ under this reparametrization are
$\begin{split}s^{*}_{i}\mid\bm{\theta}&\sim N(\alpha_{0}+\bm{w}^{\top}_{i}\bm{\alpha},1)\,,\\ s^{*}_{i}\mid\{y^{*}_{i},\bm{\theta}\}&\sim N(\alpha_{0}+\bm{w}^{\top}_{i}\bm{\alpha}+\frac{\tilde{\rho}}{\tilde{\sigma}^{2}+\tilde{\rho}^{2}}(y_{i}-\beta_{0}-\bm{x}^{\top}_{i}\bm{\beta}),\frac{\tilde{\sigma}^{2}}{\tilde{\sigma}^{2}+\tilde{\rho}^{2}})\,,\\ y^{*}_{i}\mid\{s^{*}_{i},\bm{\theta}\}&\sim N\left(\beta_{0}+\bm{x}^{\top}_{i}\bm{\beta}+\tilde{\rho}(s^{*}_{i}-\alpha_{0}-\bm{w}_{i}^{\top}\bm{\alpha}\right),\tilde{\sigma}^{2})\,.\end{split}$ (5)
where $\bm{\theta}=(\beta_{0},\bm{\beta}^{\top},\alpha_{0},\bm{\alpha}^{\top},\tilde{\rho},\tilde{\sigma}^{2})^{\top}$. Full details about the Gibbs sampler, and the corresponding priors, are presented by van Hasselt (2011) and reproduced in Appendix C for completeness. The major advantage of this Gibbs sampler is that every conditional distribution is in closed-form, allowing for fast simulations. The sampler does rely on $(p+1)\times(p+1)$ and $(q+1)\times(q+1)$ matrix inversions but is rare for the problems of interest in sample selection literature to exhibit dimensions high enough for this to pose a serious concern. The Gibbs sampler also requires sampling from $n$ truncated normal random variables at each step, but this can be done relatively fast using available methods and R packages (Geweke, 1991). We will capitalize on van Hasselt (2011) to develop a Gibbs sampler in Section 4 for the priors proposed in Section 3.
3 Spike-and-slab prior formulation
Our aim is to perform Bayesian variable selection in sample selection models of type (1). To this end, let us define the variable inclusion indicators as follows. For $j=1,\ldots,p$, let $\gamma^{O}_{j}=1$ if $\beta_{j}$ is included in the outcome model, and $\gamma^{O}_{j}=0$ otherwise. Similarly, for $k=1,\ldots,q$, let $\gamma^{S}_{k}=1$ if $\alpha_{k}$ is included in the selection model, and $\gamma^{S}_{k}=0$ otherwise. Next, we present the two proposed classes of priors. The first class of priors contains continuous spike-and-slab priors with scale mixture of normal components (George and McCulloch, 1993). The second class of priors is defined conditionally on the marginal variance of the outcome model. Finally, we present a discussion on elicitation for each class of priors.
3.1 Priors
The Class I spike-and-slab prior (George and McCulloch, 1993) by the structure:
$\begin{split}\bm{\beta}\mid\{\bm{\gamma^{O}},\bm{v^{O}}\}&\sim\prod_{j=1}^{p}\left((1-\gamma^{O}_{j})N(0,\tau_{0,\beta}^{2}v^{O}_{j})+\gamma_{j}^{O}N(0,\tau_{1,\beta}^{2}v^{O}_{j})\right)\,\\ \bm{\alpha}\mid\{\bm{\gamma^{S}},\bm{v^{S}}\}&\sim\prod_{k=1}^{q}\left((1-\gamma^{S}_{k})N(0,\tau_{0,\alpha}^{2}v^{S}_{k})+\gamma_{k}^{S}N(0,\tau_{1,\alpha}^{2}v^{S}_{k})\right)\,,\end{split}$ (6)
where $v_{j}^{O}$ and $v_{k}^{S}$ are positive random variables, allowing for scale mixtures of normal distributions as priors, such as Laplace and Student-$t$ distributions. Formally, we let
$p(v_{j}^{O}\mid\gamma_{j}^{O})=(1-\gamma_{j}^{O})\pi_{0,\beta}(v_{j}^{O})+\gamma_{j}^{O}\pi_{1,\beta}(v_{j}^{O})\,,$ (7)
where $\pi_{0,\beta}$ and $\pi_{1,\beta}$ are probability density functions with positive support. The distribution of $v_{k}^{S}\mid\gamma_{k}^{S}$ is defined analogously. Note that conditional on $\gamma_{j}^{O}$, only one of the components in (7) is active, so marginalizing (6) over $\bm{v^{O}}$ and $\bm{v^{S}}$ results in the components being different scale mixtures of normals, depending on the choices of $\pi_{0,\beta},\pi_{1,\beta},\pi_{0,\alpha}$ and $\pi_{1,\alpha}$. Hence, this formulation allows for different distributions for the spike and slab components. For instance, one could use a normal distribution for the spike and a Laplace distribution for the slab.
The Class II spike-and-slab prior is defined by the structure:
$\begin{split}\bm{\beta}\mid\{\bm{\gamma^{O}},\bm{v^{O}},\tilde{\sigma}^{2}\}&\sim\prod_{j=1}^{p}\left((1-\gamma^{O}_{j})N(0,\tau_{0,\beta}^{2}v^{O}_{j}\tilde{\sigma}^{2})+\gamma_{j}^{O}N(0,\tau_{1,\beta}^{2}v^{O}_{j}\tilde{\sigma}^{2})\right)\,,\\ \bm{\alpha}\mid\{\bm{\gamma^{S}},\bm{v^{S}}\}&\sim\prod_{k=1}^{q}\left((1-\gamma^{S}_{k})N(0,\tau_{0,\alpha}^{2}v^{S}_{k})+\gamma_{k}^{S}N(0,\tau_{1,\alpha}^{2}v^{S}_{k})\right)\,,\end{split}$ (8)
with $\bm{v^{O}}$ and $\bm{v^{S}}$ defined as before. In sample selection models, the variance of the observed outcomes is $\tilde{\sigma}^{2}$, as shown in equation (5). Choosing the marginal variance $\sigma^{2}=\tilde{\sigma}^{2}+\tilde{\rho}^{2}$ would not lead to a natural expression for the posterior of $\tilde{\sigma}^{2}$, whereas $\tilde{\sigma}^{2}=\sigma^{2}(1-\rho^{2})$ does.
The advantage of Class II priors over Class I is that the magnitude of variables is only considered relative to the unexplained variance. This is a common strategy in the standard linear regression context (Louzada et al., 2023). The cost of using $\tilde{\sigma}^{2}$ instead of $\sigma^{2}$ is that the observed variance is less than the true variance, with $\tilde{\sigma}^{2}$ being shrunk dependent on $\rho$. For small magnitudes of $\rho$, the difference is small, but for large values, this can have a significant impact on the priors. For instance, when $\rho=0.7$, $\tilde{\sigma}^{2}=0.49\sigma^{2}$.
The priors for the remaining parameters, common to both classes of priors, are specified as follows. We let $\gamma^{O}_{j}$ and $\gamma^{S}_{k}$ follow i.i.d. Bernoulli distributions with parameter $r$. We let $r\sim Beta(a_{0},b_{0})$, so that the priors on $\bm{\gamma}^{O}$ and $\bm{\gamma}^{S}$ are Beta-Binomial. One could also consider different $r$’s for the outcome and selection models, aiming at separating the penalties on the corresponding model sizes. We do not pursue this option in this work as this scenario is of less interest in practice. We further assume $\tilde{\rho}\mid\tilde{\sigma}\sim N(0,\tau\tilde{\sigma}^{2})$, and $\tilde{\sigma}^{2}\sim\text{IG}(c,d)$, where IG denotes an inverse gamma distribution. For the intercepts, we use weakly informative priors $\beta_{0}\sim N(0,\eta^{O}v^{O}_{0})$ and $\alpha_{0}\sim N(0,\eta^{S}v^{S}_{0})$, with large values of $\eta^{S}$ and $\eta^{O}$, and $v^{O}_{0}\sim\pi_{I,\beta}$ and $v^{S}_{0}\sim\pi_{I,\alpha}$ to allow for scale mixtures of normals as priors for the intercepts.
3.2 Prior calibration
Posterior inference is sensitive to the choice of the spike and slab variances (Tadesse and Vannucci, 2021). In this section we discuss intuitive choices that we found to work reasonably well in practice. The performance of this strategy will be illustrated through a simulation study in Section 5.
With $r\sim Beta(a_{0},b_{0})$, a natural choice is $a_{0}=b_{0}=1$, which is a uniform prior on $(0,1)$. This choice is reasonable in sparse scenarios, but it allows for $\mathbb{E}(r\mid\bm{\gamma^{O}},\bm{\gamma^{S}})>0.5$, that is the prior inclusion probability to be greater than $0.5$. An alternative is using $a_{0}=1,b_{0}=p+q$, so that $\mathbb{E}(r\mid\bm{\gamma^{O}},\bm{\gamma^{S}})\leq 0.5$ in every case except the full model. The trade-off is the increased penalty as the dimension increases. Continuous spike-and-slab priors are sensitive to the choice of variances, so we give particular importance to their calibration. We calibrate the variances for the normal prior case, as calibration for scale mixtures of normals will depend on the scale distribution and their tail-weight. While theoretical results are not available for sample selection models, we take inspiration from Narisetty and He (2014) and impose $\tau_{0,\alpha}\rightarrow 0$ as $n\rightarrow\infty$ (and similarly for $\tau_{0,\beta}$). In sparse situations, the Beta-Binomial prior imposes greater bias against small effects as $p$ and $q$ increase. This is because each $\gamma^{O}_{j}$ and $\gamma^{S}_{k}$ are drawn from a Bernoulli with the parameter $r$ having a $Beta(1,1)$ hyperprior over it. The posterior for $r$ depends on the number of variables included - the less variables included, the smaller the expected value of $r\mid\{\bm{\gamma^{O}},\bm{\gamma^{S}}\}$. For very sparse situations, $r$ is shrunk to small values, so that the prior inclusion probability becomes very small. As a result, the Beta-Binomial prior avoids the multiple testing problem when there are many irrelevant variables, but for small sample sizes it can lead to the exclusion of small effects more often than desired. To partially offset this, we propose the choice $\tau_{0,\alpha}=1/\sqrt{nc_{p}}$ and $\tau_{0,\beta}=1/\sqrt{nc_{q}}$. The factor $1/\sqrt{n}$ is motivated by theoretical results for linear regression models for spike-and-slab priors (Narisetty and He, 2014; Narisetty et al., 2019). The constants $c_{p}$ and $c_{q}$ may depend on $p$ and $q$. This does not entirely offset the increasing bias against small effects, however. For smaller sample sizes, the prior variance is still influential in the posterior. While smaller spike variances make it less likely for coefficient samples to be excluded, if they are excluded by chance, later samples in the chain will be shrunk to small values, so that it takes many iterations for the small effect to be included again. Regardless, we recommend $c_{p}=p$ and $c_{q}=q$ as heuristic choices, as we found these choices to partially offset this bias and make the inference slightly less dependent on dimension.
We similarly choose $\tau_{1,\alpha}$ and $\tau_{1,\beta}$ in line with previous work on spike-and-slab priors. For the probit equation, we choose $\tau_{1,\alpha}=C_{\alpha}$, a constant. This is in line with the choice made by Narisetty et al. (2019) for logistic regression. We choose $C_{\alpha}=\sqrt{3}/\pi$ to match the probit scale. For the linear part of the model, we use $\tau_{1,\beta}=C_{\beta}(\log(n))^{1/2}$, which was used by Narisetty and He (2014) in their simulations. The discussion in Section 3.4 of Tadesse and Vannucci (2021) suggests that for this choice, as $n\rightarrow\infty$, the probability of including a non-zero effect will converge to $1$. This would be true for a fixed choice $\tau_{1,\beta}$ not dependent on $n$ as well, but the elicitation of $\tau_{1,\beta}$ is dependent on sample size. Even for moderately large sample sizes, a small choice of $\tau_{1,\beta}$ can shrink large coefficients and negatively affect the inference. But for small sample sizes, large $\tau_{1,\beta}$ will exclude small effects, losing sensitivity. Since this is not as much of an issue for moderately large sample sizes, we suggest a slab variance proportional to $\log(n)$.
The elicitation of $C_{\beta}$ still depends on the context, and our default choice is derived as follows. Suppose that $\tau_{0,\beta}=1/\sqrt{n}$ and assume that $r=0.5$. For a small effect of magnitude $0.25$, there is variability in the posterior sampling, so we wish to ensure that we include marginally smaller effects, say of magnitude $0.15$, with high probability, but not effects significantly smaller. With $\tau_{1,\beta}=0.5$, the probability of including a coefficient sample of magnitude $0.15$ is approximately $96\%$. But an effect size of $0.05$ only has a $14\%$ probability of inclusion. As a result, we recommend the heuristic choice of $C_{\beta}=\left(4\log(500)\right)^{-1/2}$. It should be noted that while the above recommendations provide a reasonable trade-off between including small effects and excluding spurious ones, it may be desirable to focus on one over the other. For instance, if there is no interest in including small effects, one could use a larger value for $C_{\beta}$ such as $1$ to improve specificity.
We do not directly define priors on $\rho$ and $\sigma^{2}$, but indirectly through $\tilde{\rho}$ and $\tilde{\sigma}^{2}$. The prior on $\tilde{\sigma}^{2}$ is $\text{IG}(c,d)$, with $c,d>0$ positive constants, and the prior on $\tilde{\rho}$ is $N(0,\tau\tilde{\sigma}^{2})$ where $\tau>0$ is a positive constant. We let the prior on $\tilde{\rho}$ depend on $\tilde{\sigma}^{2}$ so that the induced prior on $\rho$ has its density approach zero as $\rho\rightarrow\pm 1$. Choosing $c=d=1$ leads to an induced prior on $\sigma^{2}$ with similar behaviour to an Inverse Gamma. The choice of $\tau$ in the prior for $\tilde{\rho}$ has little impact on the posterior inference for $\bm{\gamma^{S}},\bm{\gamma^{O}},\alpha_{0},\beta_{0},\bm{\alpha},\bm{\beta}$ and $\sigma^{2}$. However, it does have an impact on the posterior inference of $\rho$. The induced prior on $\rho$ depends on the choice of $\tau$: for small values such as $\tau=0.5$, it behaves similarly to a uniform distribution, while for larger values like $\tau=5$, it behaves similarly to a $Beta(1/2,1/2)$ distribution. Empirically we found that $\tau=5$ leads to more accurate posterior means of $\rho$ than $\tau=0.5$. See Appendix B for a further discussion.
4 Gibbs samplers
In this section, we formulate Gibbs samplers for the sample selection model (1) coupled with the priors structures proposed in Section 3. After initializing all model parameters,
  • Step 1.
    For $i=1,\ldots,n$, sample from $\bm{s^{*}}$, where $s^{*}_{i}\mid\{y_{i},s_{i},\alpha_{0},\beta_{0},\bm{\alpha},\bm{\beta},\tilde{\rho},\tilde{\sigma}\}$ is distributed according to
    $\begin{cases}\mathcal{TN}_{(-\infty,0)}(\alpha_{0}+\bm{w}_{i}^{\top}\bm{\alpha},1),&\text{if }y_{i}\text{ is missing},\\ \mathcal{TN}_{(0,\infty)}\left(\alpha_{0}+\bm{w}_{i}^{\top}\bm{\alpha}+\frac{\tilde{\rho}}{\tilde{\rho}^{2}+\tilde{\sigma}^{2}}(y_{i}-\beta_{0}-\bm{x}_{i}^{\top}\bm{\beta}),\frac{\tilde{\sigma}^{2}}{\tilde{\rho}^{2}+\tilde{\sigma}^{2}}\right),&\text{otherwise},\end{cases}$
    where $\mathcal{TN}_{(a,b)}$ denotes a truncated normal distribution on $(a,b)$.
  • Step 2.
    Sample $(\alpha_{0},\bm{\alpha}^{\top})^{\top}$ from $N(\bm{a^{*}},\bm{A^{*}})$.
  • Step 3.
    Jointly sample $(\beta_{0},\bm{\beta}^{\top},\tilde{\rho})^{\top}$ from $N(\bm{b^{*}},\bm{B^{*}})$.
  • Step 4.
    Sample $\tilde{\sigma}^{2}$ from $\text{IG}(c^{*},d^{*})$, where
    $c^{*}$ $=c+\frac{1}{2}\left(1+\sum_{i=1}^{n}s_{i}\right)\,,$
    $d^{*}$ $=d+\frac{\tilde{\rho}^{2}}{2\tau}+\frac{1}{2}\sum_{\{i:s_{i}=1\}}\left(y_{i}-\beta_{0}-\bm{x}_{i}^{\top}\bm{\beta}-\tilde{\rho}(s^{*}_{i}-\alpha_{0}-\bm{w}_{i}^{\top}\bm{\alpha})\right)^{2}\,.$
  • Step 5.
    For $j=1,\ldots,p$, sample $\gamma^{O}_{j}\mid\{\beta_{j},v^{O}_{j}\}\sim Ber(r^{O}_{j})$ where
    $r^{O}_{j}=\frac{r\phi\left(\beta_{j}/\left(\tau_{1,\beta}\sqrt{v^{O}_{j}}\right)\right)\pi_{1,\beta}(v^{O}_{j})}{r\phi\left(\beta_{j}/\left(\tau_{1,\beta}\sqrt{v^{O}_{j}}\right)\right)\pi_{1,\beta}(v^{O}_{j})+(1-r)\phi\left(\beta_{j}/\left(\tau_{0,\beta}\sqrt{v^{O}_{j}}\right)\right)\pi_{0,\beta}(v^{O}_{j})}\,.$
  • Step 6.
    For $k=1,\ldots,q$, sample $\gamma^{S}_{k}\mid\{\alpha_{k},v^{S}_{k}\}\sim Ber(r^{S}_{k})$ where
    $r^{S}_{k}=\frac{r\phi\left(\alpha_{k}/\left(\tau_{1,\alpha}\sqrt{v^{S}_{k}}\right)\right)\pi_{1,\alpha}(v^{S}_{k})}{r\phi\left(\alpha_{k}/\left(\tau_{1,\alpha}\sqrt{v^{S}_{k}}\right)\right)\pi_{1,\alpha}(v^{S}_{k})+(1-r)\phi\left(\alpha_{k}/\left(\tau_{0,\alpha}\sqrt{v^{S}_{k}}\right)\right)\pi_{0,\alpha}(v^{S}_{k})}\,.$
  • Step 7.
    Sample $r\mid\{\bm{\gamma^{O}},\bm{\gamma^{S}}\}\sim Beta(a_{1},b_{1})$, where
    $a_{1}$ $=a_{0}+\sum_{j=1}^{p}\gamma^{O}_{j}+\sum_{k=1}^{q}\gamma^{S}_{k}\,,$
    $b_{1}$ $=b_{0}+p+q-\sum_{j=1}^{p}\gamma^{O}_{j}-\sum_{k=1}^{q}\gamma^{S}_{k}\,.$
  • Step 8.
    For $j=1,\ldots,p$, sample $v^{O}_{j}\mid\{\gamma^{O}_{j},\beta_{j}\}$ from
    $p(v^{O}_{j}\mid\gamma^{O}_{j},\beta_{j})\propto(1-\gamma^{O}_{j})\phi\left(\frac{\beta_{j}}{\tau_{0,\beta}\sqrt{v^{O}_{j}}}\right)\pi_{0,\beta}(v^{O}_{j})+\gamma^{O}_{j}\phi\left(\frac{\beta_{j}}{\tau_{1,\beta}\sqrt{v^{O}_{j}}}\right)\pi_{1,\beta}(v^{O}_{j})\,,$
    where $\pi_{0,\beta}$ and $\pi_{1,\beta}$ are the mixing distributions.
  • Step 9.
    For $k=1,\ldots,q$, sample $v^{S}_{k}\mid\{\gamma^{S}_{k},\alpha_{k}\}$ from
    $p(v^{S}_{k}\mid\gamma^{S}_{k},\alpha_{k})\propto(1-\gamma^{S}_{k})\phi\left(\frac{\alpha_{k}}{\tau_{0,\alpha}\sqrt{v^{S}_{k}}}\right)\pi_{0,\alpha}(v^{S}_{k})+\gamma^{S}_{k}\phi\left(\frac{\alpha_{k}}{\tau_{1,\alpha}\sqrt{v^{S}_{k}}}\right)\pi_{1,\alpha}(v^{S}_{k})\,;$
    where $\pi_{0,\alpha}$ and $\pi_{1,\alpha}$ are the mixing distributions.
  • Step 10.
    Sample $v^{O}_{0}\sim\phi(\beta_{0}/(\sqrt{\eta^{O}v^{O}_{0}}))\pi_{I,\beta}(v^{O}_{0})$ and $v^{S}_{0}\sim\phi(\alpha_{0}/(\sqrt{\eta^{S}v^{S}_{0}}))\pi_{I,\alpha}(v^{S}_{0})$.
The exact expressions for $\bm{a^{*}},\bm{A^{*}},\bm{b^{*}}$ and $\bm{B^{*}}$ can be found in Appendix C. Steps 1 through 7 have distributions that are well-known and can easily be sampled from. Steps 8, 9 and 10 depend on the choice of scale distributions $\pi_{0,\alpha},\pi_{1,\alpha},\pi_{0,\beta}$ and $\pi_{1,\beta}$. There is no closed form expression for a general scale mixing distribution, but certain choices lead to closed form sampling for these steps. A Laplace prior can be used by choosing the mixing distribution to be $\text{Exp}(1/2)$: in this case, the posterior distribution to sample from takes the form of an Inverse Gaussian (after conditioning in Step 8/9). A Student-$t$ prior can be used by choosing the mixing distribution to be Inverse Gamma, which is conjugate to the normal likelihood. In these cases and the normal case (by choosing the mixing distributions to be point masses at $1$), the sampler is in closed form.
The output of the above procedure will be a sample from the posterior $(\bm{\alpha}^{\top},\bm{\beta}^{\top},{\bm{\gamma^{S}}}^{\top},{\bm{\gamma^{O}}}^{\top},\tilde{\sigma}^{2},\tilde{\rho})^{\top}\mid\{\bm{y},\bm{s}\}$. While the sampler still relies on $(p+1)\times(p+1)$ and $(q+1)\times(q+1)$ matrix inversions, the dimensions are usually not so big in applications of interest that this is problematic.
5 Simulation study
In this section, we present a simulation study that aims at illustrating how the prior calibration performs as the dimension increases in highly sparse situations, for small and moderately large samples and typical levels of missingness in the data. We compare the performance of the Gibbs samplers to alternatives such as Adaptive LASSO and stepwise selection.
5.1 Simulation scenarios
The true model is of the form in (1) with $\bm{\alpha}=\left(0.5,1,1.5,0,\ldots,0\right)/\sqrt{2}\in\mathbb{R}^{q}$, $\bm{\beta}=\left(0.25,0.5,1,0,\ldots,0\right)\in\mathbb{R}^{p}$, that is $3$ active variables and $p-3$ spurious variables, and $p=q$. The effect sizes are chosen to represent a small, medium large effect respectively. The effect sizes for the probit equation are in line with Certo et al. (2016). The intercept $\alpha_{0}$ is chosen dependent on the simulated covariates, so that the expected proportion of missing data is $0.3$. We choose $\beta_{0}=0.5$ and $\sigma=1$. We generated $n$ covariates $\bm{w}_{i}\in\mathbb{R}^{q}$ such that their marginal distributions are standard normal and $\text{cov}(w_{j},w_{k})=0.5^{\lvert j-k\rvert}$. We let $\bm{w}_{i}=\bm{x}_{i}$ so that there is no exclusion restriction. For each $\bm{w}_{i}$, we generated values of $y_{i}$ and $s_{i}$, and repeated this for $1000$ Monte Carlo replicates (keeping the covariates fixed for each scenario). We consider sample sizes $n=500$ and $1000$, and dimensions $p=10,25$ and $50$, We performed simulations for $\rho=0,0.3,0.5$ and $0.7$, and found that the performance of each method was not dependent on $\rho$. For this reason, we only present the results for $\rho=0.5$ in the main text, and include results for other values of $\rho$ in Appendix G, along with results for $n=2000$ in Appendix I. We compare the following variable selection methods:
  1. 1.
    A normal-normal Class I spike-and-slab prior, with hyperparameters elicitated as in Section 3.2.
  2. 2.
    A Laplace-Laplace Class I spike-and-slab prior, with hyperparameters as in the normal-normal case, choosing the scale parameter of the Laplace distributions such that they have equal variances to the normal case.
  3. 3.
    Adaptive LASSO as in Ogundimu (2022).
  4. 4.
    A vanilla implementation of forward selection which starts from the null model and, at each step, fits the maximum likelihood estimates of the parameters and adds the variable which minimizes the Akaike Information Criterion (AIC) to either the selection or outcome equation (whichever minimizes the AIC).
5.2 Performance measures
The evaluation metrics we use are: Sensitivity (TPR) - the proportion of active variables correctly identified; Specificity (TNR) - the proportion of inactive variables correctly identified; True model rate (TMR) - the proportion of replicates where the selected model is the true model; Model size - the average size of the selected model. For spike-and-slab samplers, the “selected model” is the median model, that is the model such that a variable is included if and only if the posterior inclusion probability (PIP) is greater than $0.5$. Formally, a variable $\beta_{j}$ is “included” if more than half of the posterior samples of $\gamma^{O}_{j}$ are equal to $1$, and similarly for $\alpha_{k}$. Barbieri and Berger (2004) shows in the normal linear case that this is the optimal predictive model.
We run the Gibbs samplers for $10,000$ iterations, discarding the first $1,250$ iterations as burn-in. These choices provide a compromise between convergence of the chain and computational efficiency, and increasing the length of the chain has diminishing returns. The sampleSelection R package is used to choose initial values for the spike-and-slab samplers. That is, the parameter estimates from the maximum likelihood fit are used as the starting parameters, and the initial value of $\bm{\gamma}^{S}$ and $\bm{\gamma}^{O}$ are chosen such that a variable is included in the model if and only if its parameter estimate (under the maximum likelihood fit) is significant at a $5\%$ level. For both the normal and Laplace cases, we choose $\tau=5$ for the prior on $\tilde{\rho}$, $Beta(1,1)$ for the prior on $r$, and use the same distribution for the intercept as the rest of the model, with the chosen slab variance as its variance. For the normal case, we let $\tau_{1,\alpha}=\tau_{1,\beta}=0.5$, and we divide these by $\sqrt{2}$ for the Laplace case.
It should be noted that for some cases, commonly used maximum likelihood estimators fail to converge, causing stepwise selection and Adaptive LASSO to fail. To mitigate this, we ran additional replicates until we had $1000$ replicates that had converged, and report the proportion of seeds that lead to failed convergence. The proposed Gibbs samplers do not have this problem, so we do not simulate additional replicates. The chosen initial values do depend on this, so in the event that the maximum likelihood estimators fail to converge, the initial values of all the parameters are set to 00 except $\tilde{\sigma}=1$ and $r=0.5$, and the initial model is set to the null model.
5.3 Results
The results of the simulations can be found in Appendix F. Table 7 shows the performance of each method for $n=500$ and $p=10,25$ and $50$ respectively. It should be noted that the medium and large effects are included with probability close to one, so that the sensitivity is almost entirely determined by the inclusion of the small effect. The normal spike-and-slab Gibbs sampler has the highest true model rate for all choices of $p$ considered, though the Laplace spike-and-slab is close enough in this measure of performance that it cannot be said that the normal spike-and-slab estimates the true model more often overall. The Laplace spike-and-slab includes the small effect more frequently, at the cost of a higher rate of false positives. We emphasize that these results for spike-and-slab priors depend on prior elicitation. For example, a larger slab variance for either prior distribution would improve specificity but exclude the small effect more often.
As the sparsity of the true model increases, the performance of every method deteriorates. The methods based on spike-and-slab priors exclude the small effect more often with increasing sparsity; this is a result of the multiple testing penalty imposed by the Beta-Binomial prior. Adaptive LASSO in particular struggles as the sparsity increases. It increasingly includes irrelevant variables in the selection equation - averaging a model size of $4.57$ for $p=50$ when the true model size is $3$ and increasingly excludes the small effect in the outcome equation, to a larger degree than either of the spike-and-slab priors do. Stepwise selection performs poorly in every situation by all metrics considered, particularly so as the dimension increases.
Table 8 compares the methods for $n=1000$. All methods perform significantly better, with Adaptive LASSO producing a competitive performance at this sample size. Adaptive LASSO exhibits better performance for the outcome equation in low dimensions, while the spike-and-slab priors outperform Adaptive LASSO in the presence of sparsity when taking the selection equation into account. Stepwise selection still performs worse than the other methods.
When $n=500$ and both $p$ and $\rho$ are large, both Adaptive LASSO and stepwise selection suffer from convergence issues. For $p=50$, a large proportion of the replicates failed. For $\rho=0$, 37.2% cases failed to converge; for $\rho=0.3$, 51.8% cases failed; for $\rho=0.5$, 71.3% cases failed; and for $\rho=0.7$, 91.8% cases failed. Stepwise selection suffers from a similar problem to a lesser extent, with 1.28%, 1.28%, 4.85% and 13.9% of the replicates failing for the same values of $\rho$. This presents a serious problem, as irrespective of the performance of the methods, there is a high chance the algorithms fail to converge. The Gibbs samplers formulated in this paper do not exhibit this issue and can be expected to work well when other methods fail to converge. In addition, discarding these problematic replicates induces a bias in the results that favours Adaptive LASSO, as these replicates were still used for the spike-and-slab samplers. Even without taking that into consideration, the spike-and-slab samplers remain competitive. It should be noted that issues with convergence only happens for small samples: the problem effectively disappears for $n=1,000$. These issues could also be related to the absence of any exclusion restriction in the true model.
6 Real data applications
In this section, we present two real data applications that illustrate the use of the proposed Bayesian variable selection methodology and how it compares against adaptive LASSO and stepwise selection. The first example analyzes ambulatory expenditures (Cameron and Trivedi, 2010), whereas the second example analyzes data from the RAND Health Insurance Experiment (RAND HIE), a comprehensive study on the effect of health insurance on medical expenditures. For these applications, we standardize all the covariates and examine the parameter estimates and standard deviations using the standardized data, as the true models are unknown. For the spike-and-slab methods, we additionally look at the posterior inclusion probabilities and use the same priors as in the simulation studies (Section 5), apart from using $N(0,100)$ priors on $\alpha_{0}$ and $\beta_{0}$ (to allow for large intercepts), a $Beta(1,p+q)$ prior on $r$ (to account for low sparsity) and letting $\tau_{1,\alpha}=\sqrt{3}/\pi$ and $\tau_{1,\beta}=\left(4\log(500)\right)^{-1/2}(\log(n))^{1/2}$. The samplers are run for $50,000$ iterations, with the first $5,000$ being burn-in. We report the posterior mean of $\rho$, based on the proposed spike-and-slab priors, as a Bayes’ estimate. We additionally report the results from fitting the full model using the sampleSelection R package. We present the results for Class II priors and additional results in the Appendix. We briefly consider post-selection inference. To compare models without re-fitting the data, we follow the approach of Liang et al. (2008) by comparing parameter posteriors conditional on the given models, i.e. $\left(\bm{\alpha},\bm{\beta},\sigma,\rho\right)\mid\bm{\gamma}$. In our case we consider the sub-sample where the sampled model is $\bm{\gamma}$ (and only considering parameters with $\gamma_{j}=1$). We evaluate the performance of different models using the Bayesian leave-one-out estimate of the expected log pointwise predictive density, referred to as $elpd_{loo}$ (Vehtari et al., 2017). We use the loo package in R to compute this.
6.1 Ambulatory expenditures data
The ambulatory expenditures data contains information about several explanatory variables, such as age, gender, education status (educ), ethnicity (blhisp), number of chronic conditions (totchr), insurance status (ins) and income. The main aim is to study the effect of these variables on ambulatory expenditures. Not all patients had money spent, so that there are missing outcomes, and since we expect the decision to spend to be linked with the cost, this is a case of sample selection bias. The dataset contains $n=3,328$ observations, of which $516$ ($15.8\%$) have missing expenditure. This dataset has been studied in previous publications (Marchenko and Genton, 2012; Ogundimu, 2022), and in line with these references we use log-expenditure (lambexp) as the outcome variable. We choose the same predictors as in previous publications, for comparison. That is, we let $\bm{x}=(\texttt{age},\texttt{female},\texttt{educ},\texttt{blhisp},\texttt{totchr},\texttt{ins})^{\top}$ and $\bm{w}=(\bm{x}^{\top},\texttt{income})^{\top}$ so that income is an exclusion restriction. The use of the income for this purpose is questionable (Cameron and Trivedi, 2010), and an advantage of using variable selection methods in this context is to determine whether such restrictions are necessary.
Table 1 shows the results of applying each method to the data. The variables chosen by the spike-and-slab normal prior and the Adaptive LASSO are similar. Both exclude educ and ins from the outcome equation, with posterior inclusion probabilities being small for both variables. Stepwise selection excludes ins but includes educ. The major difference is in the exclusion restriction. The Adaptive LASSO still includes the exclusion restriction income, albeit shrunk more than the other included variables. The spike-and-slab normal, on the other hand, excludes income from the selection equation, with posterior inclusion probability $0.349$. It is still included in a considerable number of posterior samples, so that the posterior mean is not shrunk to near-zero, unlike ins and educ in the outcome equation, though it is still shrunk significantly more than Adaptive LASSO. The variable ins is also somewhat shrunk in the selection equation, with inclusion probability $0.571$. The larger standard deviations in each case is due to the bimodality of the posterior when the inclusion probability is not close to 00 or $1$.
Spike-and-slab normal ALASSO Stepwise Full model
PIP Est. S.D. Est. S.D. Est. S.D. Est. S.D.
Selection equation
(Intercept) - 1.276 0.038 1.270 0.038 1.283 0.038 1.283 0.038
educ 1.000 0.176 0.032 0.158 0.030 0.159 0.031 0.159 0.031
age 0.949 0.108 0.039 0.089 0.030 0.099 0.031 0.099 0.031
income 0.349 0.029 0.042 0.054 0.034 0.072 0.035 0.072 0.035
female 1.000 0.320 0.031 0.320 0.030 0.331 0.030 0.331 0.030
totchr 1.000 0.603 0.055 0.609 0.055 0.615 0.055 0.615 0.055
blhisp 1.000 -0.171 0.029 -0.161 0.028 -0.168 0.029 -0.168 0.029
ins 0.571 0.049 0.046 0.062 0.030 0.081 0.030 0.082 0.030
Outcome equation
(Intercept) - 6.564 0.060 6.585 0.054 6.508 0.053 6.514 0.055
educ 0.116 0.008 0.018 0.000 0.000 0.049 0.027 0.048 0.027
age 1.000 0.230 0.026 0.222 0.026 0.237 0.026 0.238 0.026
female 1.000 0.158 0.032 0.141 0.030 0.177 0.029 0.174 0.030
totchr 1.000 0.399 0.031 0.390 0.030 0.420 0.030 0.417 0.030
blhisp 0.895 -0.089 0.039 -0.076 0.028 -0.102 0.027 -0.101 0.028
ins 0.033 -0.002 0.008 0.000 0.000 0.000 0.000 -0.014 0.025
$\sigma$ 1.288 0.024 1.281 0.024 1.270 0.018 1.271 0.018
$\rho$0.265 0.154 -0.318 0.139 -0.114 0.141 -0.131 0.147
Table 1: Results from ambulatory data. ALASSO refers to Adaptive LASSO, Stepwise refers to forward selection and “Full model” to the full model fit using the sampleSelection package. “PIP” refers to posterior inclusion probabilities for each variable, “Est.” refers to parameter estimates (posterior means for spike-and-slab) and “S.D.” refers to standard deviations of the parameter estimates.
Another difference is in the estimation of $\rho$. The full model estimates a significantly smaller value of $\rho$ compared to the Adaptive LASSO estimates and posterior mean of the spike-and-slab normal. The spike-and-slab normal priors and Adaptive LASSO produce estimates of $\rho$ further away from zero.
For post-selection inference, we compared the $10$ most sampled models, which included the median model for spike-and-slab and the model returned by Adaptive LASSO, with results reported in 2. The best model and Adaptive LASSO include more variables than spike-and-slab, but the gain in predictive performance is relatively small, suggesting these additional variables contribute only modestly to predictive power.
$elpd_{loo}$ Model size
Best model -5086.0 12
Adaptive LASSO model -5087.5 11
Spike-and-slab median model -5087.7 10
Table 2: Values of $elpd_{loo}$ for the ten most sampled models under spike-and-slab for the ambulatory data. The $elpd_{loo}$ for each model was computed using only parameter samples corresponding to iterations where the given model was sampled. The “best model” refers to the model with the lowest $elpd_{loo}$ of the considered models.
6.2 RAND data
We now analyze a data set from the RAND Health Insurance Experiment (RAND HIE). This is a comprehensive study from $1971$ to $1986$ of the impact of randomized health insurance on health care cost, utilization and outcomes. This data set was used by Cameron and Trivedi (2005) to analyze how the patient’s use of health services is affected by types of randomly assigned health insurance. In line with previous applications, we use the variable lnmeddol, the logarithm of medical expenses per individual, as the outcome variable. We let $\bm{x}$ consist of the logarithm of coinsurance rate plus 1 ($\texttt{logc}=\log(\texttt{coins}+1)$), the dummy variable for individual deductible plan (idp), the logarithm of participation incentive payment (lpi), an artificial variable fmde that is 00 if $\texttt{idp}=1$ and $\log(\max(1,\texttt{mde}/(0.01*\texttt{coins}))$ otherwise (where mde is the maximum expenditure offer), physical limitations (physlm), the number of chronic diseases (disea), dummy variables for good (hlthg), fair (hlthf) and poor (hlthp) self-rated health (where the baseline is excellent self-rated health), the log of family income (linc), the log of family size (lfam), education of household head in years (educdec), age of individual in years (xage), a dummy variable for female individuals (female), a dummy variable for individuals younger than $18$ years (child), a dummy variable for female individuals younger than $18$ years (fchild), and a dummy variable for black household heads (black). The selection variable is binexp which indicates whether the medical expenses are positive, and let $\bm{w}=\bm{x}$ so that there is no exclusion restriction. A subsample is selected so that the study year is $2$ and educdec is not “NA”. There are $5,574$ observations, of which $1,293$ have zero medical expenses so that lnmeddol is missing. The table in Appendix 6 contains the results of each method applied to the subsample. All the methods estimate similar values for $\rho$ of around $0.73$, aside from stepwise selection which estimates $\rho=-0.03$, failing to detect any significant sample selection bias. Stepwise selection also excludes logc from the outcome, which both Adaptive LASSO and the spike-and-slab sampler include as a significant variable. The inference is otherwise similar between the normal spike-and-slab prior and Adaptive LASSO. In the selection equation, the same variables are excluded. The only difference in this regard is that the normal spike-and-slab prior excludes hlthg from the outcome, with the inclusion probability being $0.217$, whereas the Adaptive LASSO does not exclude this variable, though it does heavily shrink it to almost the same extent as the normal spike-and-slab prior. We also observe that the normal spike-and-slab prior produces larger estimates of the coefficients than Adaptive LASSO does.
Figure 1: Visualisation of a 50% credible set for the model space posterior. The black bars above variables indicate groups - if a line is not broken between two variables, they are considered to be in the same group. The numbers indicate whether a variable is included or not and the color indicates the posterior inclusion probability of each listed group permutation.
We take advantage of the recent advancements in Griffin (2024) to visualise the posterior on the model space. Figure 1 shows a 50% credible set for the model space posterior. Of most interest from this visualisation is the relation of the group consisting of lfam, xage and child, in particular the latter two. xage and child have a correlation of $-0.804$, and from Figure 1, it is clear that there is significant overlap in the information each variable brings into the model, as all three combinations involving one (or both) have similar posterior inclusion probability. In this case, both variables have inclusion probabilities greater than $0.5$, but if a higher cutoff were to be used, the visualisation suggests that at least one of these variables should be included regardless. It is also worth noting that hlthg and hlthf are grouped by this visualisation, which is where the only difference between the Adaptive LASSO model and the median model of the spike-and-slab occurs.
For post-selection inference, we consider the 20 most sampled models and use the spike-and-slab samples to obtain $elpd_{loo}$ values as in table 3. The best model in this case coincides with the model chosen by Adaptive LASSO, which includes one additional variable, but the gain in predictive performance by including this variable is modest.
$elpd_{loo}$ Model size
Adaptive LASSO model (best model) -8541.6 23
Spike-and-slab median model -8548.2 22
Table 3: Values of $elpd_{loo}$ for the 20 most sampled models under spike-and-slab for the RAND data. The $elpd_{loo}$ for each model was computed using only parameter samples corresponding to iterations where the given model was sampled. The “best model” refers to the model with the lowest $elpd_{loo}$ of the considered models.
6.3 Sensitivity analysis
In this section, we perform a sensitivity analysis for the spike and slab variances on the two datasets.
Let $\tau_{1,\alpha}^{*}=\sqrt{3}/\pi$ and $\tau_{1,\beta}^{*}=(4\log(500))^{-1/2}(\log(n))^{1/2}$. Alongside those values, which are the defaults for the previous analysis, we consider two additional specifications. We consider $\tau_{1,\alpha}=2\tau_{1,\alpha}^{*}$ and $\tau_{1,\beta}=2\tau_{1,\beta}^{*}$, alongside the choices $\tau_{1,\alpha}=0.5\tau_{1,\alpha}^{*}$ and $\tau_{1,\beta}=0.5\tau_{1,\beta}^{*}$. In each case, the variances have been multiplied or shrunk by a factor of $4$, so that the values have change significantly. Other than the choice of spike-and-slab variances, we use the same elicitation as in the previous two sections. The table 21 in Appendix J shows the comparison between the three different choices on the ambulatory data. Similar tables for the spike variance and analogous tables for the RAND data can also be found in the table in Appendix J. As expected, the larger slab variances have smaller inclusion probabilities for every variable. Despite the difference in prior variances, though, the difference between the models is not drastic. The only difference in the maximum a posteriori model is the exclusion of ins in the selection equation by the larger slab variances. The posterior mean of the coefficients for variables with inclusion probability not close to $1$ are shrunk more by the larger variances, but otherwise are similar. Similar behaviour occurs for the RAND data (the table in 23). Due to the high collinearity between the predictors, some variables actually have higher inclusion probability under the larger slab variance, when highly correlated variables are excluded in their place. The tables in 22 and 24 show that posterior inference based on the spike-and-slab prior is less sensitive to the spike variance than the slab variance. There is little difference in inclusion probability for most variables.
7 Conclusion
We proposed two continuous spike-and-slab prior structures for variable selection in the context of sample selection models, and provided practical guidelines for calibrating these priors. We developed Gibbs samplers, with closed-form tractable conditionals, for sample selection models coupled with the proposed prior structures. An appealing feature of the proposed prior structures and Gibbs samplers is that they allow for using (different) scale mixture of normals for the spike and slab components. We have shown that these Gibbs samplers are scalable to the dimensions of interest in practice. The time complexity is not linear, as clearly seen in Section Appendix A, so it will not scale well to dimensions $p>>200$. However, applications of interest studied in sample selection literature are seldom higher dimensional than this.
Our simulation study shows that the proposed Bayesian variable selection methodology offers a good performance in terms of sensitivity and specificity of the selected variables. Moreover, it exhibits similar performance to Adaptive LASSO in low dimensions and large sample sizes, while identifying the true model more frequently for extremely sparse models and smaller sample sizes. A significant advantage of our proposed methodology is that, unlike Adaptive LASSO, our proposed methodology does not suffer from convergence issues for moderately small sample sizes and higher dimensions.
There exist several natural extensions of our work. It is desirable to extend theoretical results, similar to Narisetty and He (2014), to sample selection models, and to use these results to inform prior elicitation, or use a data-driven approach to elicitate the hyperparameters instead (Tadesse and Vannucci, 2021). Regarding scalability, one could extend the Skinny Gibbs algorithm (Narisetty et al., 2019) to sample selection, which would let the algorithm scale to higher dimensions, or to develop ad hoc variational Bayes methods (Tadesse and Vannucci, 2021). A similar approach could be taken with EM algorithms, such as the spike-and-slab LASSO by Ročková and George (2018). We assumed bivariate normal errors in the two-equation model (1). It would be possible to extend this model to non-normal errors, such as scale mixtures of normals (or other distributions with a tractable stochastic representation), which would allow to use the Gibbs samplers developed in Section 4 with additional steps. The proposed Gibbs samplers can also be extended to the case with binary outcomes, by using the stochastic representation of probit models through latent variables. The focus in this paper is primarily on model selection and the problem of post-selection inference requires further investigation. Because the continuous spike-and-slab returns posterior probabilities for each model, an extension of the methodology in this paper could incorporate Bayesian model averaging (Raftery et al., 1997; Steel, 2020) into the output of the sampler directly. The issue of excluded parameters having positive posterior density away from zero under continuous spike-and-slab priors would need to be carefully considered in this context.
Acknowledgements
Adam Iqbal was supported by the Heilbronn Institute for Mathematical Research via funding from the EPSRC grant “Additional Funding Programme for Mathematical Sciences” (EP/V521917/1).
Appendix A Computational times
To test computational times for the spike-and-slab Gibbs sampler, we recorded the running time of the method on one replicate from the simulation study. That is, we simulated $\bm{x}_{i}=(x_{i,1},\ldots,x_{i,p})^{\top}$ as i.i.d. normal for $i=1,\ldots,n$, let $\bm{w}=\bm{x}$, then generated bivariate normal errors to obtain $(y_{i},s_{i})^{\top}$ from (1), with $\sigma=1$. We chose $\rho=0.5$ and to avoid convergence issues, $n=1000$. For the spike-and-slab normal, the chain was run for $10,000$ iterations, discarding the first $1,250$ as burn-in, as in the simulation study. We varied $p$ from $25$ to $200$ (and let $q=p$) to see how the running times scale to very high dimensions. The computations were performed on a computer with 16GB of RAM and an AMD Ryzen 5 3600 GPU. The reason this was run separately from the rest of the simulation study is because the original simulations were run in parallel and only considered up to $p=50$.
Method $p=25$ $p=50$ $p=100$ $p=200$
Spike-and-slab normal 15.5 39.3 130.7 609.0
Table 4: Computational times for the spike-and-slab prior, for $n=1000$. All times are in seconds.
Appendix B Induced prior for $\rho$
Recall that $\tilde{\sigma}^{2}\sim IG(c,d)$ and $(\tilde{\rho}\mid\tilde{\sigma}^{2})\sim N(0,\tau\tilde{\sigma}^{2})$, with $c,d,\tau>0$ positive constants. $\tau$ controls the behaviour of the induced prior on $\rho$. Figure 2 shows the shape of two induced priors on $(0,1)$, one with $\tau=0.5$ and the other with $\tau=5$. The case $\tau=5$ closely resembles a $Beta(1/2,1/2)$ prior, while $\tau=0.5$ resembles a uniform prior, aside from both approaching zero at the bounds 00 and $1$.
Figure 2: The induced priors on $\rho$ for $\tau=5$ and $0.5$ respectively. Smaller choices of $\tau$ are closer to a uniform distribution, while larger choices resemble Beta distributions, except approaching zero at the tails.
Alongside the simulation study in Section 5, we ran the same simulations for the normal-normal spike-and-slab sampler with $\tau=0.5$ (and otherwise the same choices of prior parameters). Table 5 shows how the posterior mean of $\rho$ differed between these two priors. Both priors underestimate $\rho$, but the prior with $\tau=5$ consistently gives estimates closer to the true value. The posterior inference for other parameters (not shown here) is similar for both choices of $\tau$, so we prefer $\tau=5$ to $0.5$.
n True $\rho$ $\tau=0.5$ $\tau=5$
500 0 -0.059 -0.067
500 0.3 0.174 0.193
500 0.5 0.351 0.390
500 0.7 0.559 0.616
1000 0 -0.014 -0.015
1000 0.3 0.245 0.260
1000 0.5 0.437 0.459
1000 0.7 0.642 0.669
2000 0 -0.002 -0.002
2000 0.3 0.277 0.285
2000 0.5 0.476 0.488
2000 0.7 0.676 0.690
Table 5: Comparison of posterior mean of $\rho$ for $\tau=0.5$ and $\tau=5$. Uses spike-and-slab normal priors for each case, with the same elicitation and same $1000$ replicates as in the simulation study. The posterior means of $\rho$ are very similar for all $p$, so we only include $p=50$ for brevity.
Appendix C Gibbs sampler for sample selection models
Here, we reproduce the Gibbs sampler from [van Hasselt, 2011] for completeness. Let the priors on the parameters be
$(\alpha_{0},\bm{\alpha}^{\top})^{\top}$ $\sim N(\bm{a},\bm{A})\,,$
$(\beta_{0},\bm{\beta}^{\top})^{\top}$ $\sim N(\bm{b},\bm{B})\,,$
$\tilde{\rho}\mid\tilde{\sigma}$ $\sim N(0,\tau\tilde{\sigma}^{2})\,,$
$\tilde{\sigma}^{2}$ $\sim\text{IG}(c,d)\,.$
where $\bm{a}\in\mathbb{R}^{p+1}$, $\bm{A}\in\mathbb{R}^{(p+1)\times(p+1)}$, $\bm{b}\in\mathbb{R}^{q+1}$, $\bm{B}\in\mathbb{R}^{(q+1)\times(q+1)}$, $\tau\in(0,1)$ and $c,d>0$ are real constants.
Further define $\bm{X}=(\bm{1},\bm{x}_{1},\ldots,\bm{x}_{n})$ and $\bm{W}=(\bm{1},\bm{w}_{1},\ldots,\bm{w}_{n})$. Let $\bm{W}_{0}$ be the design matrix $\bm{W}$ with only the observations that have missing $y_{i}$, and $\bm{W}_{1}$ the design matrix with observations that have non-missing $y_{i}$. Similarly, let $\bm{X}_{1}$ be the design matrix $\bm{X}$ with the observations that have non-missing $y_{i}$, $\bm{y}_{1}$ be the non-missing outcomes, $\bm{s^{*}}_{1}$ the corresponding values of $s^{*}$ and $\bm{s^{*}}_{0}$ the values of $s^{*}$ for observations with missing outcomes. Furthermore, let $\bm{g}=(\bm{\beta}^{\top},0)^{\top}$ and
$\bm{G}=\begin{pmatrix}\bm{B}&\bm{0}\\ \bm{0}&\tau\tilde{\sigma}^{2}\end{pmatrix}\,,$
and let $\bm{Z}$ be the design matrix $\bm{X}_{1}$ with an additional column $\bm{s^{*}}_{1}-\bm{W}_{1}\bm{\alpha}$. Then the Gibbs sampler as in van Hasselt [2011] is as follows:
  1. 1.
    Sample from $\bm{s^{*}}$, where for $i=1,\ldots,n$ where $s^{*}_{i}\mid\{y_{i},s_{i},\alpha_{0},\beta_{0},\bm{\alpha},\bm{\beta},\tilde{\rho},\tilde{\sigma}\}$ is distributed according to
    $\begin{cases}\mathcal{TN}_{(-\infty,0)}(\alpha_{0}+\bm{w}_{i}^{\top}\bm{\alpha},1)&\text{if }y_{i}\text{ is missing},\\ \mathcal{TN}_{(0,\infty)}\left(\alpha_{0}+\bm{w}_{i}^{\top}\bm{\alpha}+\frac{\tilde{\rho}}{\tilde{\rho}^{2}+\tilde{\sigma}^{2}}(y_{i}-\beta_{0}-\bm{x}_{i}^{\top}\bm{\beta}),\frac{\tilde{\sigma}^{2}}{\tilde{\rho}^{2}+\tilde{\sigma}^{2}}\right)&\text{otherwise}\end{cases}$
    where $\mathcal{TN}_{(a,b)}$ denotes a truncated normal distribution on $(a,b)$.
  2. 2.
    Sample $(\alpha_{0},\bm{\alpha}^{\top})^{\top}$ from $N(\bm{a^{*}},\bm{A^{*}})$, where
    $\bm{A^{*}}$ $=\left(\bm{A}^{-1}+\bm{W}_{0}^{\top}\bm{W}_{0}+\left(\frac{\tilde{\sigma}^{2}}{\tilde{\sigma}^{2}+\tilde{\rho}^{2}}\right)^{-1}\bm{W}_{1}^{\top}\bm{W}_{1}\right)^{-1}\,,$
    $\bm{a^{*}}$ $=\bm{A^{*}}\left(\bm{A}^{-1}\bm{a}+\bm{W}_{0}^{\top}\bm{s^{*}}_{0}+\left(\frac{\tilde{\sigma}^{2}}{\tilde{\sigma}^{2}+\tilde{\rho}^{2}}\right)^{-1}\bm{W}_{1}^{\top}\left(\bm{s^{*}}_{1}-\frac{\tilde{\rho}}{\tilde{\sigma}^{2}+\tilde{\rho}^{2}}(\bm{y}_{1}-\bm{X}_{1}\bm{\beta})\right)\right)\,.$
  3. 3.
    Jointly sample $(\beta_{0},\bm{\beta}^{\top},\tilde{\rho})^{\top}$ from $N(\bm{b^{*}},\bm{B^{*}})$, where
    $\bm{B^{*}}$ $=\left(\bm{G}^{-1}+\frac{1}{\tilde{\sigma}^{2}}\bm{Z}^{\top}\bm{Z}\right)^{-1}\,,$
    $\bm{b^{*}}$ $=\bm{B^{*}}\left(\bm{G}^{-1}\bm{g}+\frac{1}{\tilde{\sigma}^{2}}\bm{Z}^{\top}\bm{y}_{1}\right)\,.$
  4. 4.
    Sample $\tilde{\sigma}^{2}$ from $\text{IG}(c^{*},d^{*})$, where
    $c^{*}$ $=c+\frac{1}{2}\left(1+\sum_{i=1}^{n}s_{i}\right)\,,$
    $d^{*}$ $=d+\frac{\tilde{\rho}^{2}}{2\tau}+\frac{1}{2}\sum_{\{i:s_{i}=1\}}\left(y_{i}-\beta_{0}-\bm{x}_{i}{{}^{\top}}\bm{\beta}-\tilde{\rho}(s^{*}_{i}-\alpha_{0}-\bm{w}_{i}^{\top}\bm{\alpha})\right)^{2}\,.$
Appendix D Further details on the Gibbs samplers
D.1 Class I prior
The prior variance matrix for the outcome equation, $\bm{B}$, corresponds to the $(p+1)\times(p+1)$ diagonal matrix where $B_{1,1}=\eta^{O}v^{O}_{0}$ and $B_{j,j}=\left((1-\gamma^{O}_{j-1})\tau_{0,\beta}^{2}v+\gamma^{O}_{j-1}\tau_{1,\beta}^{2}\right)v^{O}_{j-1}$ (with the mixing variable $v^{O}_{j-1}$ appearing to allow for scale mixture of normal priors). $\bm{A}$ is defined similarly. Define $\bm{W}_{0},\bm{W}_{1},\bm{s^{*}}_{0},\bm{s^{*}}_{1},\bm{y}_{1},\bm{g},\bm{G}$ and $\bm{Z}$ as in Section C. Then we have that
$\bm{A^{*}}$ $=\left(\bm{A}^{-1}+\bm{W}_{0}^{\top}\bm{W}_{0}+\left(\frac{\tilde{\sigma}^{2}}{\tilde{\sigma}^{2}+\tilde{\rho}^{2}}\right)^{-1}\bm{W}_{1}^{\top}\bm{W}_{1}\right)^{-1}\,,$
$\bm{a^{*}}$ $=\bm{A^{*}}\left(\bm{W}_{0}^{\top}\bm{s^{*}}_{0}+\left(\frac{\tilde{\sigma}^{2}}{\tilde{\sigma}^{2}+\tilde{\rho}^{2}}\right)^{-1}\bm{W}_{1}^{\top}\left(\bm{s^{*}}_{1}-\frac{\tilde{\rho}}{\tilde{\sigma}^{2}+\tilde{\rho}^{2}}(\bm{y}-\bm{X}\bm{\beta})\right)\right)\,,$
$\bm{B^{*}}$ $=\left(\bm{G}^{-1}+\frac{1}{\tilde{\sigma}^{2}}\bm{Z}^{\top}\bm{Z}\right)^{-1}\,,$
$\bm{b^{*}}$ $=\bm{B^{*}}\left(\frac{1}{\tilde{\sigma}^{2}}\bm{Z}^{\top}\bm{y}_{1}\right)\,.$
D.2 Class II prior
The Gibbs sampler for the Class II prior is similar to the Class I prior, with only two differences. The first difference is that wherever $\tau_{0,\beta}$ and $\tau_{1,\beta}$ appear, $\tau_{0,\beta}\tilde{\sigma}$ and $\tau_{1,\beta}\tilde{\sigma}$ appear instead. The second difference are the values of $c^{*}$ and $d^{*}$ in sampling $\tilde{\sigma}^{2}$ (in Step 4 of the algorithm). That is, we instead have that
$c^{*}$ $=c+\frac{1}{2}\left(1+\sum_{i=1}^{n}s_{i}\right)+\frac{p+1}{2}\,,$
$d^{*}$ $=d+\frac{\tilde{\rho}^{2}}{2\tau}+\frac{1}{2}\sum_{\{i:s_{i}=1\}}\left(y_{i}-\bm{x}_{i}{{}^{\top}}\bm{\beta}-\tilde{\rho}(s^{*}_{i}-\bm{w}_{i}^{\top}\bm{\alpha})^{2}\right)+\frac{\beta_{0}^{2}}{2\eta^{O}v^{O}_{0}}$
$+\sum_{j=1}^{p}\frac{\beta_{j}^{2}}{2v^{O}_{j}((1-\gamma^{O}_{j})\tau_{0,\beta}^{2}+\gamma^{O}_{j}\tau_{1,\beta}^{2})}\,.$
Appendix E RAND data table
Spike-and-slab normal ALASSO Stepwise Full model
PIP Est. S.D. Est. S.D. Est. S.D. Est. S.D.
Selection equation
(Intercept) - 0.814 0.020 0.801 0.020 0.822 0.020 0.815 0.020
logc 1.000 -0.230 0.030 -0.198 0.022 -0.218 0.023 -0.218 0.054
idp 0.094 -0.004 0.012 0.000 0.000 -0.061 0.020 -0.048 0.022
lpi 0.919 0.068 0.028 0.043 0.020 0.077 0.022 0.079 0.023
fmde 0.090 0.006 0.023 0.000 0.000 0.000 0.000 0.003 0.055
physlm 0.998 0.099 0.024 0.073 0.022 0.089 0.024 0.092 0.023
disea 1.000 0.154 0.023 0.150 0.023 0.158 0.024 0.143 0.024
hlthg 0.020 0.000 0.004 0.000 0.000 0.000 0.000 0.028 0.021
hlthf 0.131 0.007 0.018 0.000 0.000 0.046 0.022 0.060 0.022
hlthp 0.844 0.070 0.038 0.053 0.023 0.076 0.026 0.099 0.025
linc 0.640 0.041 0.034 0.024 0.020 0.061 0.020 0.068 0.020
lfam 0.019 0.000 0.005 0.000 0.000 0.000 0.000 -0.017 0.022
educdec 0.922 0.068 0.027 0.050 0.019 0.102 0.021 0.089 0.021
xage 0.031 -0.001 0.007 0.000 0.000 0.000 0.000 -0.010 0.035
female 1.000 0.196 0.024 0.164 0.023 0.223 0.027 0.205 0.027
child 0.057 0.002 0.011 0.000 0.000 0.041 0.027 0.026 0.039
fchild 1.000 -0.143 0.024 -0.111 0.022 -0.179 0.032 -0.157 0.031
black 1.000 -0.227 0.021 -0.223 0.019 -0.235 0.020 -0.225 0.020
Outcome equation
(Intercept) - 3.551 0.039 3.561 0.038 4.013 0.096 3.543 0.036
logc 1.000 -0.229 0.027 -0.209 0.024 0.000 0.000 -0.155 0.069
idp 0.033 -0.001 0.007 0.000 0.000 -0.040 0.024 -0.066 0.029
lpi 0.046 0.002 0.010 0.000 0.000 0.000 0.000 0.040 0.028
fmde 0.037 -0.001 0.011 0.000 0.000 -0.129 0.025 -0.081 0.067
physlm 1.000 0.125 0.025 0.102 0.024 0.083 0.023 0.115 0.024
disea 1.000 0.205 0.027 0.204 0.026 0.133 0.027 0.194 0.026
hlthg 0.236 0.014 0.026 0.026 0.023 0.070 0.023 0.075 0.025
hlthf 0.740 0.058 0.040 0.060 0.023 0.097 0.023 0.120 0.026
hlthp 0.964 0.097 0.031 0.086 0.023 0.097 0.021 0.124 0.023
linc 0.997 0.137 0.034 0.108 0.027 0.114 0.028 0.148 0.028
lfam 0.853 -0.072 0.037 -0.060 0.024 -0.074 0.025 -0.086 0.027
educdec 0.076 0.004 0.014 0.000 0.000 0.000 0.000 0.050 0.026
xage 0.749 0.102 0.071 0.100 0.037 0.095 0.037 0.096 0.041
female 1.000 0.279 0.035 0.252 0.031 0.168 0.035 0.275 0.032
child 0.670 -0.100 0.083 -0.109 0.043 -0.131 0.044 -0.097 0.048
fchild 1.000 -0.234 0.044 -0.204 0.037 -0.136 0.039 -0.224 0.039
black 1.000 -0.205 0.030 -0.197 0.029 -0.072 0.039 -0.207 0.029
$\sigma$ 1.571 0.029 1.561 0.029 1.393 0.015 1.570 0.028
$\rho$ 0.725 0.039 0.721 0.038 -0.030 0.174 0.736 0.034
Table 6: Results from RAND health data. ALASSO refers to Adaptive LASSO, Stepwise refers to forward selection and “Full model” to the full model fit using the sampleSelection package.
Appendix F Main simulation tables
F.1 $n=500$
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.783 3.176 0.986 0.969 0.808 3.066 0.976 0.980
Laplace 0.645 3.426 0.990 0.935 0.726 3.240 0.983 0.959
Stepwise 0.355 4.002 0.999 0.856 0.363 3.986 0.996 0.857
ALASSO 0.558 3.458 0.982 0.927 0.762 2.919 0.945 0.988
$p=25$
Normal 0.764 3.043 0.964 0.993 0.794 2.964 0.957 0.996
Laplace 0.735 3.128 0.970 0.990 0.775 3.035 0.964 0.994
Stepwise 0.046 6.269 0.997 0.851 0.036 6.159 0.999 0.856
ALASSO 0.420 3.900 0.980 0.956 0.702 2.874 0.927 0.996
$p=50$
Normal 0.741 2.923 0.941 0.998 0.703 2.833 0.920 0.998
Laplace 0.720 3.003 0.947 0.997 0.698 2.914 0.931 0.997
Stepwise 0.001 10.295 0.998 0.845 0.002 10.194 0.997 0.847
ALASSO 0.271 4.570 0.973 0.965 0.537 2.717 0.871 0.998
Table 7: Evaluation metrics for $n=500,\rho=0.5$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
F.2 $n=1000$
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.876 3.138 1.000 0.980 0.929 3.079 1.000 0.989
Laplace 0.799 3.243 1.000 0.965 0.863 3.157 1.000 0.978
Stepwise 0.350 4.013 1.000 0.855 0.360 4.004 1.000 0.857
ALASSO 0.877 3.123 0.999 0.982 0.974 2.980 0.992 1.000
$p=25$
Normal 0.908 3.094 0.999 0.996 0.928 3.066 0.999 0.997
Laplace 0.871 3.145 0.999 0.993 0.896 3.105 0.999 0.995
Stepwise 0.048 6.077 1.000 0.860 0.044 6.018 1.000 0.863
ALASSO 0.650 3.493 1.000 0.978 0.938 3.033 0.995 0.998
$p=50$
Normal 0.919 3.077 0.999 0.998 0.943 3.041 0.997 0.999
Laplace 0.876 3.127 0.999 0.997 0.918 3.079 0.998 0.998
Stepwise 0.002 9.855 1.000 0.854 0.001 9.625 1.000 0.859
ALASSO 0.476 3.833 0.999 0.982 0.938 3.009 0.991 0.999
Table 8: Evaluation metrics for $n=1000,\rho=0.5$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Appendix G Additional tables for different correlations
G.1 $n=500$
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.791 3.165 0.986 0.970 0.803 3.045 0.970 0.981
Laplace 0.663 3.425 0.993 0.936 0.723 3.237 0.983 0.959
Stepwise 0.361 3.983 0.998 0.859 0.356 3.972 0.996 0.859
ALASSO 0.561 3.535 0.991 0.920 0.753 2.941 0.947 0.986
$p=25$
Normal 0.786 3.050 0.969 0.993 0.776 2.948 0.952 0.996
Laplace 0.747 3.138 0.974 0.990 0.765 3.055 0.967 0.993
Stepwise 0.037 6.259 0.997 0.852 0.034 6.194 0.996 0.854
ALASSO 0.388 4.010 0.986 0.952 0.691 2.890 0.928 0.995
$p=50$
Normal 0.739 3.005 0.954 0.997 0.701 2.868 0.924 0.998
Laplace 0.712 3.076 0.958 0.996 0.711 2.956 0.940 0.997
Stepwise 0.001 10.401 0.997 0.842 0.001 10.030 0.997 0.850
ALASSO 0.242 4.704 0.982 0.963 0.643 2.865 0.912 0.997
Table 9: Evaluation metrics for $n=500,\rho=0$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.764 3.217 0.988 0.964 0.794 3.053 0.971 0.980
Laplace 0.650 3.473 0.993 0.929 0.708 3.266 0.983 0.955
Stepwise 0.352 4.025 0.997 0.852 0.346 4.006 0.995 0.854
ALASSO 0.554 3.509 0.984 0.920 0.764 2.941 0.948 0.986
$p=25$
Normal 0.763 3.049 0.965 0.993 0.782 2.957 0.953 0.996
Laplace 0.726 3.147 0.972 0.989 0.762 3.066 0.966 0.992
Stepwise 0.035 6.311 0.997 0.849 0.043 6.166 0.998 0.856
ALASSO 0.404 3.980 0.983 0.953 0.690 2.880 0.924 0.995
$p=50$
Normal 0.731 2.944 0.941 0.997 0.708 2.847 0.922 0.998
Laplace 0.714 3.017 0.948 0.996 0.716 2.920 0.935 0.998
Stepwise 0.002 10.265 0.998 0.845 0.002 10.046 0.996 0.850
ALASSO 0.264 4.609 0.976 0.964 0.567 2.750 0.881 0.998
Table 10: Evaluation metrics for $n=500,\rho=0.3$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.795 3.156 0.984 0.971 0.832 3.045 0.977 0.984
Laplace 0.693 3.384 0.990 0.941 0.760 3.214 0.985 0.963
Stepwise 0.366 4.021 0.998 0.853 0.340 4.035 0.998 0.851
ALASSO 0.612 3.343 0.977 0.941 0.779 2.905 0.946 0.991
$p=25$
Normal 0.773 3.005 0.960 0.994 0.796 2.966 0.959 0.996
Laplace 0.722 3.107 0.965 0.990 0.783 3.034 0.967 0.994
Stepwise 0.031 6.245 0.997 0.852 0.038 6.176 0.998 0.855
ALASSO 0.426 3.788 0.973 0.961 0.713 2.895 0.932 0.996
$p=50$
Normal 0.727 2.939 0.941 0.998 0.715 2.839 0.923 0.999
Laplace 0.711 3.002 0.947 0.997 0.729 2.908 0.935 0.998
Stepwise 0.001 9.926 0.998 0.852 0.001 9.909 0.998 0.853
ALASSO 0.223 4.689 0.965 0.962 0.515 2.667 0.859 0.998
Table 11: Evaluation metrics for $n=500,\rho=0.7$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
G.2 $n=1000$
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.881 3.124 0.999 0.982 0.925 3.085 1.000 0.988
Laplace 0.799 3.230 0.999 0.967 0.858 3.166 1.000 0.976
Stepwise 0.367 3.977 1.000 0.860 0.362 3.956 1.000 0.863
ALASSO 0.876 3.122 0.999 0.982 0.959 2.971 0.988 0.999
$p=25$
Normal 0.905 3.097 0.998 0.995 0.926 3.057 0.997 0.997
Laplace 0.855 3.162 0.999 0.992 0.896 3.095 0.997 0.995
Stepwise 0.031 6.211 1.000 0.854 0.047 6.013 1.000 0.863
ALASSO 0.582 3.580 0.999 0.974 0.927 3.026 0.992 0.998
$p=50$
Normal 0.916 3.078 0.998 0.998 0.922 3.066 0.997 0.998
Laplace 0.891 3.103 0.998 0.998 0.891 3.106 0.998 0.998
Stepwise 0.001 9.854 1.000 0.854 0.003 9.537 1.000 0.861
ALASSO 0.443 3.937 1.000 0.980 0.930 3.028 0.992 0.999
Table 12: Evaluation metrics for $n=1000,\rho=0$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.880 3.125 0.999 0.982 0.924 3.081 1.000 0.988
Laplace 0.785 3.253 1.000 0.964 0.870 3.147 1.000 0.979
Stepwise 0.353 3.987 1.000 0.859 0.348 4.044 1.000 0.851
ALASSO 0.853 3.146 0.998 0.978 0.973 2.983 0.993 0.999
$p=25$
Normal 0.911 3.085 0.999 0.996 0.921 3.069 0.998 0.997
Laplace 0.868 3.140 0.999 0.994 0.883 3.117 0.999 0.995
Stepwise 0.029 6.067 1.000 0.861 0.042 6.018 1.000 0.863
ALASSO 0.619 3.519 1.000 0.976 0.929 3.035 0.993 0.998
$p=50$
Normal 0.901 3.090 0.998 0.998 0.941 3.044 0.996 0.999
Laplace 0.873 3.123 0.998 0.997 0.913 3.078 0.997 0.998
Stepwise 0.002 9.844 1.000 0.854 0.000 9.711 1.000 0.857
ALASSO 0.458 3.906 1.000 0.981 0.940 3.017 0.992 0.999
Table 13: Evaluation metrics for $n=1000,\rho=0.3$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.894 3.125 1.000 0.982 0.931 3.071 1.000 0.990
Laplace 0.814 3.234 1.000 0.967 0.866 3.149 1.000 0.979
Stepwise 0.358 3.980 1.000 0.860 0.352 4.037 1.000 0.852
ALASSO 0.915 3.080 0.998 0.988 0.984 2.986 0.995 1.000
$p=25$
Normal 0.909 3.090 0.999 0.996 0.938 3.054 0.999 0.997
Laplace 0.873 3.138 0.999 0.994 0.911 3.087 0.999 0.996
Stepwise 0.055 6.070 1.000 0.860 0.047 6.046 1.000 0.862
ALASSO 0.667 3.449 0.999 0.980 0.945 3.035 0.997 0.998
$p=50$
Normal 0.911 3.087 0.999 0.998 0.935 3.053 0.997 0.999
Laplace 0.875 3.132 1.000 0.997 0.904 3.092 0.998 0.998
Stepwise 0.001 9.834 1.000 0.855 0.000 9.707 1.000 0.857
ALASSO 0.477 3.847 0.999 0.982 0.934 3.029 0.993 0.999
Table 14: Evaluation metrics for $n=1000,\rho=0.7$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Appendix H Class II prior comparison
Recall that $\tilde{\sigma}^{2}=\sigma^{2}(1-\rho^{2})$. For $\rho=0$, the Class II prior is exactly dependent on the outcome variance, and the prior is shrunk for $\lvert\rho\rvert>0$. In the simulation studies, we chose $\sigma^{2}=1$, so the Class II prior performs similarly to the Class I prior for small $\rho$; for this reason, we present results for the most extreme case, $\rho=0.7$.
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Class I normal 0.795 3.156 0.984 0.971 0.832 3.045 0.977 0.984
Class II normal 0.785 3.196 0.987 0.966 0.791 3.129 0.981 0.973
Class I Laplace 0.693 3.384 0.990 0.941 0.760 3.214 0.985 0.963
Class II Laplace 0.655 3.493 0.994 0.927 0.677 3.369 0.988 0.942
$p=25$
Class I normal 0.773 3.005 0.960 0.994 0.796 2.966 0.959 0.996
Class II normal 0.780 3.025 0.964 0.994 0.786 3.002 0.963 0.995
Class I Laplace 0.722 3.107 0.965 0.990 0.783 3.034 0.967 0.994
Class II Laplace 0.746 3.135 0.973 0.990 0.759 3.093 0.971 0.992
$p=50$
Class I normal 0.727 2.939 0.941 0.998 0.715 2.839 0.923 0.999
Class II normal 0.760 2.931 0.946 0.998 0.724 2.877 0.931 0.998
Class I Laplace 0.711 3.002 0.947 0.997 0.729 2.908 0.935 0.998
Class II Laplace 0.734 3.011 0.953 0.997 0.729 2.918 0.937 0.998
Table 15: Comparison between Class I and Class II priors for $n=500,\rho=0.7$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model.
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Class I normal 0.894 3.125 1.000 0.982 0.931 3.071 1.000 0.990
Class II normal 0.892 3.130 1.000 0.981 0.893 3.116 1.000 0.983
Class I Laplace 0.814 3.234 1.000 0.967 0.866 3.149 1.000 0.979
Class II Laplace 0.809 3.246 1.000 0.965 0.795 3.246 1.000 0.965
$p=25$
Class I normal 0.909 3.090 0.999 0.996 0.938 3.054 0.999 0.997
Class II normal 0.916 3.085 0.999 0.996 0.914 3.080 0.999 0.996
Class I Laplace 0.873 3.138 0.999 0.994 0.911 3.087 0.999 0.996
Class II Laplace 0.869 3.146 0.999 0.993 0.876 3.129 0.999 0.994
$p=50$
Class I normal 0.911 3.087 0.999 0.998 0.935 3.053 0.997 0.999
Class II normal 0.916 3.084 0.999 0.998 0.922 3.071 0.998 0.998
Class I Laplace 0.875 3.132 1.000 0.997 0.904 3.092 0.998 0.998
Class II Laplace 0.874 3.132 0.999 0.997 0.890 3.111 0.998 0.998
Table 16: Comparison between Class I and Class II priors for $n=1000,\rho=0.7$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model.
The two perform fairly similarly, with Class II priors including more variables on average than Class I priors (reflected in the model size).
Appendix I Tables for $n=2000$
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.934 3.073 1.000 0.990 0.952 3.052 1.000 0.993
Laplace 0.872 3.146 1.000 0.979 0.918 3.091 1.000 0.987
Stepwise 0.351 4.014 1.000 0.855 0.381 3.926 1.000 0.868
ALASSO 0.998 3.000 1.000 1.000 0.901 2.901 0.967 1.000
$p=25$
Normal 0.932 3.072 1.000 0.997 0.963 3.037 1.000 0.998
Laplace 0.897 3.115 1.000 0.995 0.941 3.060 1.000 0.997
Stepwise 0.055 6.100 1.000 0.859 0.038 6.140 1.000 0.857
ALASSO 0.945 3.057 1.000 0.997 1.000 3.000 1.000 1.000
$p=50$
Normal 0.933 3.068 1.000 0.999 0.943 3.057 1.000 0.999
Laplace 0.892 3.116 1.000 0.998 0.925 3.076 1.000 0.998
Stepwise 0.002 9.702 1.000 0.857 0.001 9.658 1.000 0.858
ALASSO 0.726 3.337 1.000 0.993 0.990 3.010 1.000 1.000
Table 17: Evaluation metrics for $n=2000,\rho=0$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.926 3.079 1.000 0.989 0.966 3.036 1.000 0.995
Laplace 0.881 3.136 1.000 0.981 0.927 3.079 1.000 0.989
Stepwise 0.373 3.964 1.000 0.862 0.371 3.964 1.000 0.862
ALASSO 0.998 3.002 1.000 1.000 0.902 2.902 0.967 1.000
$p=25$
Normal 0.926 3.075 1.000 0.997 0.965 3.037 1.000 0.998
Laplace 0.899 3.114 1.000 0.995 0.944 3.060 1.000 0.997
Stepwise 0.045 6.117 1.000 0.858 0.041 6.021 1.000 0.863
ALASSO 0.943 3.058 1.000 0.997 0.999 2.999 1.000 1.000
$p=50$
Normal 0.938 3.064 1.000 0.999 0.969 3.031 1.000 0.999
Laplace 0.915 3.090 1.000 0.998 0.933 3.070 1.000 0.999
Stepwise 0.001 9.526 1.000 0.861 0.001 9.679 1.000 0.858
ALASSO 0.728 3.333 1.000 0.993 0.997 3.003 1.000 1.000
Table 18: Evaluation metrics for $n=2000,\rho=0.3$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.935 3.068 1.000 0.990 0.960 3.042 1.000 0.994
Laplace 0.879 3.132 1.000 0.981 0.927 3.078 1.000 0.989
Stepwise 0.359 3.993 1.000 0.858 0.363 3.974 1.000 0.861
ALASSO 1.000 3.000 1.000 1.000 0.925 2.925 0.975 1.000
$p=25$
Normal 0.937 3.071 1.000 0.997 0.965 3.037 1.000 0.998
Laplace 0.903 3.106 1.000 0.995 0.943 3.059 1.000 0.997
Stepwise 0.049 6.087 1.000 0.860 0.048 6.019 1.000 0.863
ALASSO 0.961 3.043 1.000 0.998 1.000 3.000 1.000 1.000
$p=50$
Normal 0.939 3.066 1.000 0.999 0.964 3.037 1.000 0.999
Laplace 0.914 3.096 1.000 0.998 0.943 3.060 1.000 0.999
Stepwise 0.002 9.511 1.000 0.861 0.002 9.579 1.000 0.860
ALASSO 0.754 3.297 1.000 0.994 0.997 3.004 1.000 1.000
Table 19: Evaluation metrics for $n=2000,\rho=0.5$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.941 3.066 1.000 0.991 0.963 3.039 1.000 0.994
Laplace 0.888 3.125 1.000 0.982 0.924 3.079 1.000 0.989
Stepwise 0.364 3.979 1.000 0.860 0.369 3.970 1.000 0.861
ALASSO 0.999 2.999 1.000 1.000 0.936 2.936 0.979 1.000
$p=25$
Normal 0.947 3.054 1.000 0.998 0.970 3.032 1.000 0.999
Laplace 0.916 3.090 1.000 0.996 0.940 3.064 1.000 0.997
Stepwise 0.035 6.140 1.000 0.857 0.057 5.978 1.000 0.865
ALASSO 0.975 3.025 1.000 0.999 1.000 3.000 1.000 1.000
$p=50$
Normal 0.946 3.058 1.000 0.999 0.969 3.031 1.000 0.999
Laplace 0.905 3.102 1.000 0.998 0.943 3.059 1.000 0.999
Stepwise 0.001 9.680 1.000 0.858 0.002 9.481 1.000 0.862
ALASSO 0.780 3.261 1.000 0.994 0.996 3.004 1.000 1.000
Table 20: Evaluation metrics for $n=2000,\rho=0.7$. “TMR” refers to the proportion of seeds where the true model was selected and “Size” to the number of variables in the model. Stepwise uses forward selection and ALASSO refers to Adaptive LASSO.
Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Appendix J Additional tables for sensitivity analysis
J.1 Varying slab variance for ambulatory data
Slab variance scaling ×$\times 1$$\times 0.25$$\times 4$
PIP Est. S.D. PIP Est. S.D. PIP Est. S.D.
Selection equation
(Intercept) - 1.276 0.038 - 1.269 0.037 - 1.276 0.038
educ 1.000 0.176 0.032 1.000 0.172 0.032 0.996 0.180 0.033
age 0.949 0.108 0.039 0.977 0.108 0.034 0.914 0.108 0.044
income 0.349 0.029 0.042 0.471 0.037 0.044 0.206 0.019 0.037
female 1.000 0.320 0.031 1.000 0.319 0.030 1.000 0.320 0.031
totchr 1.000 0.603 0.055 1.000 0.586 0.052 1.000 0.604 0.055
blhisp 1.000 -0.171 0.029 1.000 -0.169 0.029 1.000 -0.174 0.029
ins 0.571 0.049 0.046 0.693 0.057 0.044 0.388 0.035 0.044
Outcome equation
(Intercept) - 6.564 0.060 - 6.562 0.057 - 6.566 0.059
educ 0.116 0.008 0.018 0.196 0.011 0.021 0.062 0.005 0.014
age 1.000 0.230 0.026 1.000 0.229 0.026 1.000 0.231 0.026
female 1.000 0.158 0.032 1.000 0.158 0.030 1.000 0.157 0.031
totchr 1.000 0.399 0.031 1.000 0.398 0.030 1.000 0.398 0.031
blhisp 0.895 -0.089 0.039 0.952 -0.092 0.033 0.844 -0.085 0.042
ins 0.033 -0.002 0.008 0.079 -0.002 0.010 0.017 -0.001 0.008
$\sigma$ 1.288 0.024 - 1.287 0.023 - 1.289 0.024
$\rho$0.265 0.154 - -0.262 0.145 - -0.271 0.150
Table 21: Results from ambulatory data from varying the slab variances. The first column refers to what $\tau_{1,\alpha}$ and $\tau_{1,\beta}$ have been scaled by. The elicitation is otherwise the same as the previous data studies.
J.2 Varying spike variance for ambulatory data
Spike variances scaling ×$\times 1$$\times 0.25$$\times 4$
PIP Est. S.D. PIP Est. S.D. PIP Est. S.D.
Selection equation
(Intercept) - 1.276 0.038 - 1.278 0.038 - 1.275 0.039
educ 1.000 0.176 0.032 1.000 0.176 0.032 1.000 0.175 0.031
age 0.949 0.108 0.039 0.944 0.106 0.040 0.910 0.105 0.041
income 0.349 0.029 0.042 0.402 0.033 0.045 0.301 0.031 0.039
female 1.000 0.320 0.031 1.000 0.322 0.031 1.000 0.321 0.030
totchr 1.000 0.603 0.055 1.000 0.606 0.055 1.000 0.602 0.056
blhisp 1.000 -0.171 0.029 1.000 -0.171 0.029 1.000 -0.172 0.029
ins 0.571 0.049 0.046 0.626 0.052 0.046 0.487 0.047 0.042
Outcome equation
(Intercept) - 6.564 0.060 - 6.556 0.059 - 6.564 0.060
educ 0.116 0.008 0.018 0.131 0.007 0.019 0.103 0.013 0.019
age 1.000 0.230 0.026 1.000 0.231 0.026 1.000 0.231 0.026
female 1.000 0.158 0.032 1.000 0.161 0.031 0.999 0.157 0.032
totchr 1.000 0.399 0.031 1.000 0.401 0.031 1.000 0.398 0.032
blhisp 0.895 -0.089 0.039 0.935 -0.093 0.036 0.816 -0.084 0.040
ins 0.033 -0.002 0.008 0.036 -0.001 0.006 0.039 -0.004 0.013
$\sigma$ 1.288 0.024 - 1.286 0.023 - 1.288 0.024
$\rho$0.265 0.154 - -0.243 0.152 - -0.265 0.155
Table 22: Results from ambulatory data from varying the spike variances. The first column refers to what $\tau_{0,\alpha}$ and $\tau_{0,\beta}$ have been scaled by. The elicitation is otherwise the same as the previous data studies.
J.3 Varying slab for RAND data
Slab variance scaling ×$\times 1$$\times 0.25$$\times 4$
PIP Est. S.D. PIP Est. S.D. PIP Est. S.D.
Selection equation
(Intercept) - 0.814 0.020 - 0.813 0.020 - 0.813 0.020
logc 1.000 -0.230 0.030 1.000 -0.229 0.033 1.000 -0.231 0.029
idp 0.094 -0.004 0.012 0.206 -0.009 0.018 0.047 -0.003 0.009
lpi 0.919 0.068 0.028 0.930 0.066 0.027 0.899 0.068 0.030
fmde 0.090 0.006 0.023 0.155 0.009 0.028 0.068 0.005 0.023
physlm 0.998 0.099 0.024 0.997 0.096 0.024 1.000 0.100 0.023
disea 1.000 0.154 0.023 1.000 0.152 0.023 1.000 0.155 0.023
hlthg 0.020 0.000 0.004 0.049 0.000 0.006 0.007 0.000 0.004
hlthf 0.131 0.007 0.018 0.336 0.017 0.026 0.051 0.003 0.011
hlthp 0.844 0.070 0.038 0.960 0.082 0.030 0.724 0.060 0.042
linc 0.640 0.041 0.034 0.849 0.055 0.029 0.552 0.036 0.035
lfam 0.019 0.000 0.005 0.046 -0.001 0.006 0.009 0.000 0.004
educdec 0.922 0.068 0.027 0.961 0.071 0.024 0.833 0.061 0.032
xage 0.031 -0.001 0.007 0.072 -0.001 0.009 0.012 0.000 0.005
female 1.000 0.196 0.024 1.000 0.195 0.024 1.000 0.196 0.024
child 0.057 0.002 0.011 0.082 0.003 0.012 0.016 0.001 0.006
fchild 1.000 -0.143 0.024 1.000 -0.141 0.025 1.000 -0.143 0.023
black 1.000 -0.227 0.021 1.000 -0.223 0.020 1.000 -0.228 0.021
Outcome equation
(Intercept) - 3.551 0.039 - 3.554 0.040 - 3.546 0.039
logc 1.000 -0.229 0.027 0.994 -0.223 0.036 1.000 -0.230 0.026
idp 0.033 -0.001 0.007 0.102 -0.004 0.015 0.013 -0.001 0.005
lpi 0.046 0.002 0.010 0.085 0.003 0.012 0.022 0.001 0.007
fmde 0.037 -0.001 0.011 0.106 -0.004 0.024 0.020 -0.001 0.009
physlm 1.000 0.125 0.025 1.000 0.120 0.025 1.000 0.129 0.025
disea 1.000 0.205 0.027 1.000 0.201 0.027 1.000 0.208 0.027
hlthg 0.236 0.014 0.026 0.497 0.029 0.033 0.084 0.005 0.017
hlthf 0.740 0.058 0.040 0.898 0.078 0.037 0.507 0.038 0.040
hlthp 0.964 0.097 0.031 0.993 0.108 0.026 0.886 0.085 0.038
linc 0.997 0.137 0.034 1.000 0.146 0.031 1.000 0.130 0.034
lfam 0.853 -0.072 0.037 0.916 -0.076 0.033 0.715 -0.061 0.043
educdec 0.076 0.004 0.014 0.175 0.008 0.021 0.036 0.002 0.010
xage 0.749 0.102 0.071 0.783 0.093 0.063 0.736 0.114 0.079
female 1.000 0.279 0.035 1.000 0.271 0.034 1.000 0.285 0.035
child 0.670 -0.100 0.083 0.799 -0.112 0.073 0.544 -0.089 0.092
fchild 1.000 -0.234 0.044 1.000 -0.223 0.042 1.000 -0.242 0.044
black 1.000 -0.205 0.030 1.000 -0.203 0.030 1.000 -0.207 0.030
$\sigma$ 1.571 0.029 - 1.576 0.030 - 1.568 0.030
$\rho$ 0.725 0.039 - 0.731 0.040 - 0.721 0.041
Table 23: Results from RAND data from varying the slab variances. The first column refers to what $\tau_{1,\alpha}$ and $\tau_{1,\beta}$ have been scaled by. The elicitation is otherwise the same as the previous data studies.
J.4 Varying spike for RAND data
Spike variance scaling ×$\times 1$$\times 0.25$$\times 4$
PIP Est. S.D. PIP Est. S.D. PIP Est. S.D.
Selection equation
(Intercept) - 0.814 0.020 - 0.813 0.020 - 0.813 0.020
logc 1.000 -0.230 0.030 1.000 -0.230 0.032 1.000 -0.231 0.032
idp 0.094 -0.004 0.012 0.115 -0.004 0.013 0.082 -0.007 0.012
lpi 0.919 0.068 0.028 0.897 0.065 0.029 0.870 0.066 0.030
fmde 0.090 0.006 0.023 0.118 0.008 0.027 0.111 0.009 0.027
physlm 0.998 0.099 0.024 0.981 0.097 0.027 0.988 0.098 0.025
disea 1.000 0.154 0.023 1.000 0.153 0.023 1.000 0.154 0.024
hlthg 0.020 0.000 0.004 0.016 0.000 0.003 0.019 0.000 0.007
hlthf 0.131 0.007 0.018 0.147 0.007 0.019 0.105 0.008 0.016
hlthp 0.844 0.070 0.038 0.916 0.077 0.033 0.809 0.068 0.038
linc 0.640 0.041 0.034 0.758 0.048 0.032 0.619 0.042 0.033
lfam 0.019 0.000 0.005 0.019 0.000 0.004 0.021 0.000 0.007
educdec 0.922 0.068 0.027 0.924 0.067 0.027 0.881 0.066 0.029
xage 0.031 -0.001 0.007 0.033 -0.001 0.006 0.028 -0.001 0.008
female 1.000 0.196 0.024 1.000 0.196 0.025 1.000 0.197 0.024
child 0.057 0.002 0.011 0.054 0.002 0.011 0.046 0.003 0.011
fchild 1.000 -0.143 0.024 1.000 -0.143 0.024 1.000 -0.145 0.024
black 1.000 -0.227 0.021 1.000 -0.225 0.021 1.000 -0.227 0.021
Outcome equation
(Intercept) - 3.551 0.039 - 3.548 0.038 - 3.550 0.039
logc 1.000 -0.229 0.027 1.000 -0.230 0.027 0.998 -0.228 0.029
idp 0.033 -0.001 0.007 0.050 -0.002 0.009 0.039 -0.003 0.010
lpi 0.046 0.002 0.010 0.047 0.002 0.009 0.044 0.003 0.011
fmde 0.037 -0.001 0.011 0.045 -0.001 0.012 0.044 -0.001 0.015
physlm 1.000 0.125 0.025 0.998 0.123 0.026 0.999 0.124 0.026
disea 1.000 0.205 0.027 1.000 0.205 0.027 1.000 0.206 0.027
hlthg 0.236 0.014 0.026 0.304 0.018 0.029 0.220 0.016 0.026
hlthf 0.740 0.058 0.040 0.823 0.065 0.038 0.714 0.059 0.040
hlthp 0.964 0.097 0.031 1.000 0.104 0.025 0.968 0.098 0.030
linc 0.997 0.137 0.034 1.000 0.145 0.031 0.999 0.137 0.033
lfam 0.853 -0.072 0.037 0.924 -0.078 0.032 0.777 -0.067 0.040
educdec 0.076 0.004 0.014 0.088 0.004 0.015 0.070 0.005 0.014
xage 0.749 0.102 0.071 0.681 0.088 0.072 0.712 0.095 0.071
female 1.000 0.279 0.035 1.000 0.277 0.034 1.000 0.278 0.034
child 0.670 -0.100 0.083 0.735 -0.113 0.083 0.730 -0.111 0.081
fchild 1.000 -0.234 0.044 1.000 -0.230 0.044 1.000 -0.231 0.043
black 1.000 -0.205 0.030 1.000 -0.205 0.029 1.000 -0.206 0.030
$\sigma$ 1.571 0.029 - 1.572 0.030 - 1.572 0.029
$\rho$ 0.725 0.039 - 0.726 0.039 - 0.728 0.038
Table 24: Results from RAND data from varying the spike variances. The first column refers to what $\tau_{0,\alpha}$ and $\tau_{0,\beta}$ have been scaled by. The elicitation is otherwise the same as the previous data studies.